OpenStudy (anonymous):

4(1-x^2)=0; find for x

OpenStudy (anonymous):

distibute the 4 over the terms in parens

OpenStudy (anonymous):

4(1-x^2)=0 -divide by 4 1-x^2=0 -add x^2 1=x^2 -sqrt sqrt1=x +\-1=x

OpenStudy (anonymous):

ups, sry: x=+-1